After 15 years,Kirans are will be 4 times as old as he is now his present age is ?
Answer by Theo(13342)

x = his age now.
x + 15 = his age 15 year from now.
(x+15) = 4x
this means that in 15 years kiran's age will be 4 times what it is now.
solve for x.
x+15 = 4x
subtract x from both sides of the equation to get:
15 = 3x
divide both sides of the equation by 3 to get:
x = 5
kiran is 5 years old now.
in 15 years kiran will be 20 years old.
20/5 = 4